OpenStudy (anonymous):

What is the answer for x-9 (45+116)

OpenStudy (anonymous):

Well first you should go to the math section of this website. I will help you though. If you are trying to solve for x then you want to get x by itself by putting it into an equation. \[x-9(45+116)=0\]

OpenStudy (anonymous):

To get x by itself we need to use PEMDAS (Parenthesise, Exponents, Multiplication, Division, Addition, Subtraction). First do the math in the parenthesise. \[45+116\]

OpenStudy (anonymous):

Our new equations is\[x-9(161)=0\]

OpenStudy (anonymous):

The next step is multiplications because there are no exponents. When a number is next to parenthesise, it is telling you to multiply that number by what is inside the parenthesise. So, we need to multiply 9 by 161.\[9*161\]

OpenStudy (anonymous):

Our new equation is \[x-1449=0\]

OpenStudy (anonymous):

Now, to get x by itself we must get the -1449 to the other side of the equation. To do that we need to add 1449 to both sides of the equation.\[x-1449+1449=+1449\]

OpenStudy (anonymous):

-1449 and +1449 cancel each other out so we are left with \[x=1449\] That is the answer.